OpenStudy (anonymous):

What is the solution to the equation 9^(x + 1) = 27 ? x = 2.5 x = 0.5 x = -0.5 x = -3.5 stuck on this!

jimthompson5910 (jim_thompson5910):

9^(x + 1) = 27 (3^2)^(x + 1) = 3^3 3^(2(x+1)) = 3^3 Since the bases are equal, the exponents are equal. So 2(x+1) = 3 I'll let you continue

OpenStudy (anonymous):

i guess it's D

OpenStudy (anonymous):

no.. you need to solve for x in jim's last line...

OpenStudy (anonymous):

don't guess...

OpenStudy (anonymous):

i'm working it out

OpenStudy (anonymous):

2x+2=3

jimthompson5910 (jim_thompson5910):

so far so good

OpenStudy (anonymous):

then 2x=1

OpenStudy (anonymous):

x=1/2?

OpenStudy (anonymous):

x=.5

jimthompson5910 (jim_thompson5910):

yes, which is what in decimal form?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

gnarly